OpenStudy (is3535):

Which statement is true about the value of (5−n)(5n)? For n<0, the value of the expression is greater than 1. For n<0, the value of the expression is less than 1. For all n, the value of the expression is 0. For all n, the value of the expression is 1.

OpenStudy (is3535):

@iGreen

OpenStudy (is3535):

@welshfella

OpenStudy (is3535):

@Jaynator495

OpenStudy (is3535):

@welshfella @iGreen

OpenStudy (is3535):

plzzz help

OpenStudy (welshfella):

try plugging in appropriate values for n for example - take the first choice:- n < 0 so try n = -1:- (5 - n)(5n) (5 - 5(-1)) (5*-1) = (5+5)(-5) = -50 so is the first choice the correct one?

OpenStudy (is3535):

omg ty so much i love for that ty

OpenStudy (welshfella):

is the value of the expression greater than 1?

OpenStudy (welshfella):

yw